# SageAttention2: Efficient Attention with Thorough Outlier Smoothing and Per-thread INT4 Quantization

### v3

@@ 3 SageAttention-2
3.1 Formulation
Based on the introduction of FlashAttention and quantization presented in Section 2, we now describe the quantized attention approach we developed.
Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q},~{}~{}\overrightarrow{Q}_{m})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{Q}}}(Q),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{K}}}(K),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V},\overrightarrow{V}_{m})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\phi_{V}}}(V)$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\Delta S=\overrightarrow{Q_{m}}K^{\top}}}$ (2)
Attention: $S={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi^{-1}_{\delta_{Q}\delta_{K}}(\hat{Q}\hat{K}^{\top})+\Delta S}},~{}~{}(m^{\prime},\widetilde{P})=\tilde{\sigma}(m,S),~{}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{P},\hat{P})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{P}}}(\widetilde{P})$
$O=\mathrm{diag}\left(\exp(m^{\prime}-m)\right)O+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi^{-1}_{\delta_{P}\delta_{V}}(\hat{P}\hat{V})}}$ (3)
$\phi_{Q}$, $\phi_{K}$, $\phi_{V}$ are three transformations to obtain quantized $Q$, $K$, and $V$, which we will discuss in subsequent sections. For simplicity, we omit all superscripts and subscripts, but the matrices used in attention are still tiles, and the computation is still organized as FlashAttention described in Section 2.1. Compared to the original full-precision version, as shown in Eq. 23, SageAttention2 adds quantizers to $Q,K,P,V$ and dequantizers to the product to accelerate both Matmuls of $QK$ and $\widetilde{P}V$.
@@ 3.1 Formulation
Table 1: End-to-end metrics comparison of different quantization methods, where $Q,K$ are quantized into INT4, while $\widetilde{P},V$ stay in full precision.
Q, K Smoothing (Q+K) Llama3.1 (Lambda) $\uparrow$ Llama3.1 (WikeText) $\downarrow$ CogVideoX (vqa-a) $\uparrow$ CogVideoX (vqa-t) $\uparrow$
Full-Precision - 81.5% 6.013 77.605 75.360
INT4 Quantization 72.6% 11.698 27.114 24.670
80.8% 6.219 77.276 75.147
@@ 3.2 Smooth Q
The representative range for quantization of INT4 is notably restrictive, i.e., only $152^{4}-1=$ 15 values. This limitation significantly degrades performance when quantizing $Q,K$ into INT4 in attention. For example, using INT4 to quantize $Q,K$ can cause the perplexity of Llama3.1 on WikiText to increase by more than 90%, and the quality of videos generated by CogvideoX decrease about 3x (See Table 1). We analyze the data distribution of $Q,K$ in real models. For example, we find that Llama3.1 and CogvideoX show distinct channel-wised outliers in $Q,K$ (See Figure 4). While per-channel quantization could mitigate the quantization error caused by such outliers, this method is impracticable for $Q,K$ because quantization must be conducted along the outer axis (token dimension) of $QK^{\top}$.
To address this challenge, we propose a method to eliminate the quantization error brought by the outliers in $Q,K$:
$\overrightarrow{Q_{m}}=\mathrm{mean}(Q),~{}~{}\gamma(Q)=$ $Q-\overrightarrow{Q_{m}},~{}~{}\gamma(K)=K-\mathrm{mean}(K),$
$S=QK^{\top}=$ $\gamma(Q)\gamma(K)^{\top}+\overrightarrow{Q_{m}}\gamma(K)^{\top}$
$\phi_{Q}(Q)=\psi_{Q}\circ\gamma(Q),~{}\overrightarrow{Q_{m}}$ $,~{}~{}~{}~{}~{}\phi_{K}(K)=\psi_{K}\circ\gamma(K)$ (4)
Where $\mathrm{mean}(K)$ and $\gamma(K)$ have been discussed in SageAttention (Zhang et al., 2024), and the transformation for $K$ does not change the attention score $\widetilde{P}$. $\mathrm{mean}(Q)=\frac{1}{N}\sum_{t=1}^{N}Q[t,:]$ is a vector with a shape $1\times d$. For the transformation of $Q$, we will perform a GEMV (general matrix-vector multiplication) between $\overrightarrow{Q_{m}}$ and $K^{\top}$, i.e., $\overrightarrow{Q_{m}}K^{\top}$ as $\Delta S$. This $\Delta S$ will be added to $S$ during the attention computation. Finally, the transformation from full-precision $Q,K$ to quantized $\hat{Q},\hat{K}$ can be expressed as shown in Eqution 4, where $\psi_{Q}$, $\psi_{K}$ is two quantizers for $Q$ and $K$. In other words, full-precision $Q,K$ are substracted with the means of the channel dimension before being quantized.
@@
Figure 5 shows an example from CogvideoX of the distribution of quantized $Q$ with and without smoothing Q. We can find that with smoothing Q, the range of INT4 is utilized more uniformly and fully. Table 1 presents end-to-end metrics for different quantization methods with and without smoothing Q+K on Llama3.1 (Dubey et al., 2024) and CogvideoX (Yang et al., 2024). The results demonstrate that smoothing Q+K offers significant accuracy benefits. Also, Table 9 and Table 9 show that the order of effectiveness is smoothing Q+K $>$ smoothing Q $>$ smoothing K $>$ other baselines.

### v4

@@ 3.1 Smooth Q
First, we discuss how to accurately compute $QK^{\top}$ with INT4. The numerical range of INT4 is notably restrictive. This affects quantization due to the presence of outliers (Lin et al., 2024). Given the INT4 range [-7, +7], any element will be quantized to zero if it is more than 14 times (0.5 vs 7) smaller than the largest element in the group. Since outliers are much larger than other elements, it is quite likely that most non-outlier elements are quantized to zero, resulting in significant accuracy degradation. Therefore, to keep the quantization accurate, we need to keep the largest element small, making the magnitude of elements as uniform as possible. Such technique is called smoothing.
Here, we propose a smoothing technique inspired by SageAttention (Zhang et al., 2025). SageAttention observed that $Q,K$ for all tokens are actually highly similar, with only small variations between different tokens (see Fig. 3). We propose to smooth $K$ as SageAttention does and further smooth $Q$ by subtracting a common mean of each block:
$\gamma(Q_{i})=Q_{i}-\bar{q}_{i},~{}~{}~{}\gamma(K_{j})=K_{j}-\bar{k},$ (2)
where $\bar{q}_{i}=\mathrm{mean}(Q_{i}),\bar{k}=\mathrm{mean}(K)$ are $1\times D$ vectors, the mean is conducted along the token axis, and $\bar{q}_{i},\bar{k}$ are broadcasted to tokens in a block and a tensor for subtraction, respectively.
With the decomposion, we have $S_{ij}=Q_{i}K_{j}^{\top}=(\bar{q}_{i}+\gamma(Q_{i}))(\bar{k}+\gamma(K_{j}))^{\top}=\bar{q}_{i}\bar{k}^{\top}+\bar{q}_{i}\gamma(K_{j})^{\top}+\gamma(Q_{i})\bar{k}^{\top}+\gamma(Q_{i})\gamma(K_{j})^{\top}=\gamma(Q_{i})\gamma(K_{j})^{\top}+\Delta{S_{ij}}+b$. Here, $\Delta{S_{ij}}=\bar{q}_{i}\gamma(K_{j})^{\top}$ is an $1\times N$ vector, and $b=\bar{q}_{i}\bar{k}^{\top}+\gamma(Q_{i})\bar{k}^{\top}$ is an $N\times 1$ vector. We do not need to compute $b$ since adding a common bias to an entire row of $S$ does not affect the result after softmax. Therefore, we can accelerate $Q_{i}K_{j}^{\top}$ with INT4 by the following two stages:
(1) preprocessing: smooth $Q,K$ according to Eq. (2), apply quantization $(\delta_{Q_{i}},\hat{Q}_{i})=\psi_{Q}(\gamma(Q_{i})),(\delta_{K_{j}},\hat{K}_{j})=\psi_{K}(\gamma(K_{j}))$, and compute $\Delta{S_{ij}}=\bar{q}_{i}\gamma(K_{j})^{\top}$. The smoothing, quantization, and the GEMV (general matrix-vector multiplication) for computing $\Delta S$ can be fused into a single kernel, which scans the off-chip $Q$ and $K$ only once.
(2) attention: execute the low-precision GEMM, dequantize, and add back the vector $\Delta S$: $S_{ij}=\psi^{-1}_{\delta_{Q_{i}}\delta_{K_{j}}}(\hat{Q}_{i}\hat{K}_{j}^{\top})+\Delta S_{ij}$. These operations are all done on chip, and the dequantization and vector addition only adds a marginal overhead compared to the expensive mma operation for MM.
Importantly, $\gamma(Q_{i}),\gamma(K_{j})$ are quantized rather than $Q_{i},K_{j}$. Since the smoothed matrices are much smaller in magnitude and contain less outliers, the quantization accuracy can be significantly improved.
Remark. Classical techniques to improve the activation-weight MM, such as per-channel quantization, or SmoothQuant (Xiao et al., 2023) are not applicable here for the query-key MM in attention. Per-channel quantization is impracticable for $Q,K$ because quantization must be conducted along the outer axis (token dimension) of $QK^{\top}$ (Xiao et al., 2023). On the other hand, both $Q$ and $K$ have significant outliers, so trading the quantization accuracy between them with SmoothQuant cannot work effectively, as shown in Sec. 4.1. Here, we utilize the unique token-similarity pattern in attention to derive an accurate dedicated quantization method for $Q$ and $K$. The previous work SageAttention only smooths $K$, so it is less accurate than our method.
@@
Empirical results. Fig. 19 in Appendix A.7 shows an example from CogvideoX of the distribution of $\hat{Q}$ with and without smoothing $Q$. We can find that with smoothing $Q$, the range of INT4 is utilized more uniformly and fully. Table 5 presents end-to-end metrics for different quantization methods with and without smoothing Q+K on Llama3.1 (Dubey et al., 2024) and CogvideoX (2b) (Yang et al., 2025). The results demonstrate that smoothing Q+K offers significant accuracy benefits. Also, Table 4 and 17 show that the order of effectiveness is smoothing Q+K $>$ smoothing Q $>$ smoothing K $>$ other baselines.
@@
3.2 INT4 Per-thread Quantization
Orthogonal to smoothing, we can mitigate the problem of outliers by refining the quantization granularity so the number of affected elements by outliers becomes smaller. Although per-token offers a detailed level of granularity, it results in significant overhead during dequantization. Specifically, each GPU thread in per-token quantization must handle multiple quantization scales, leading to a high latency of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$. SageAttention uses per-block quantization, where each block $Q_{i}$ ($b_{q}$ tokens) and $K_{i}$ ($b_{k}$ tokens) have a single quantization scale. Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the high dequantization overhead. However, quantizing $Q$ and $K$ to INT4 demands a finer quantization granularity. To address this, we propose per-thread quantization, a more precise and granular approach than the per-block quantizer, also without the additional overhead of vectors dot product between $\delta_{Q}$ and $\delta_{K}$.
Specifically, each block of $Q$, i.e., $Q_{i}$, in SageAttention will be split into $c_{w}$ segments and processed by $c_{w}$ GPU warps in a GPU streaming processor (SM). We call each segment of $Q_{i}$ as $Q_{w}$, and $k_{w}=K_{j}$ since $K_{j}$ is shared among warps. Then, each warp containing 32 threads uses the mma.m16n8k64 PTX instruction (NVIDIA, ) for the $Q_{w}K_{j}^{\top}$. According to the layout requirement of this instruction, we find that $Q_{w}[8k+i]$ could share one quantization scale, and $K_{j}[8k+2i]$ and $K_{j}[8k+2i+1]$ could share one quantization scale. Such a quantization method is more fine-grained and with no additional overhead. This is because it assigns different GPU threads to distinct quantization groups based on the MMA instruction layout, with each thread performing dequantization only using a single quantization scale value. We show an example of per-thread quantization in Fig. 4. The detailed formulation is shown in Equation A.5 and Fig. 17 (Please see Appendix A.5 for more detail).
Empirical results. As shown in Table 6 and Table 15, we compare the average and the worst accuracy of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX. Results indicate the accuracy of per-thread quantization is very close to per-token and outperforms a lot more than others.
Algorithm 1 Implementation of SageAttention2.
  Preprocessing: $K=K-\mathrm{mean}(K)$, ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$. //per-channel.
  Divide $Q$ to $T_{m}={N}/{b_{q}}$ blocks $\{Q_{i}\}$; divide $K$, and $V$ to $T_{n}={N}/{b_{kv}}$ blocks $\{K_{i}\}$, $\{V_{i}\}$;
  for $i=1$ to $T_{m}$ do
    ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\bar{q}_{i}=\mathrm{mean}(Q_{i}),~{}~{}(\delta_{Q},\hat{Q}_{i})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}(Q_{i}-\bar{q}_{i})}}$ //per-thread ;
    for j in [1, $T_{n}$do
       ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K}_{j})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K_{j})$ //per-thread,     $w=\text{range}(c_{w}),st=w*c_{w}$ ;
       $S_{ij}[st:st+c_{w}]=\psi^{-1}_{\delta_{Q}\delta_{K}}(\mathrm{Matmul}(\hat{Q}_{i}[st:st+c_{w}],\hat{K}_{j}^{\top}))+\text{GEMV}(\bar{q}_{i},K_{j}^{\top})$ ; // Paralleled by $c_{w}$ warps. The $\psi^{-1}_{\delta_{Q}\delta_{K}}$ is illustrated in Fig. 4.
       $m_{ij}=\mathrm{max}(m_{i,j-1},\mathrm{rowmax}(S_{ij}))$, ${}\widetilde{P}_{ij}=\mathrm{exp}(S_{ij}-m_{ij})$, ${}l_{ij}=e^{m_{i,j-1}-m_{ij}}+\mathrm{rowsum}(\widetilde{P}_{ij})$ ;
       $O_{ij}$(FP22) = $\mathrm{Matmul}((\widetilde{P}_{ij}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
       $O_{ij}$(FP32) = $\mathrm{diag}(e^{m_{i,j-1}-m_{ij}})^{-1}O_{i,j-1}(FP32)+O_{ij}$(FP22) ;
    end for
[…]
  end for
  return $O=\{O_{i}\}$
3.3 FP8 quantization for $\tilde{P}V$
We now turn to the MM $\tilde{P}V$, where $\tilde{P}_{ij}=\exp(S_{ij}-m_{ij})$ is the unnormalized quantity according to Eq. (1). The distribution of $\tilde{P}$ is unique and differs from other activations. First, we note that $S_{ij}-m_{ij}\leq 0$, so $P_{ij}\in[0,1]$ ($\leq$ and $\in$ apply element-wisely). We find that $\tilde{P}$ often consists of many small elements, but their sum is non-negligible (e.g., 5000 elements around $10^{-4}$). In this case, we must represent small elements accurately. INT quantization is unsuitable for this proposal since it puts the quantization points evenly within the numerical range. SageAttention (Zhang et al., 2025) choose to retain $\tilde{P}$ and $V$ in FP16, and accelerate the MM by decreasing the accumulator precision. However, this strategy is only effective on very few GPUs.
We propose to quantize $\widetilde{P},V$ to FP8 with 4 exponent bits and 3 mantissa bits (E4M3). The numerical range of E4M3 is $[-448,+448]$. We quantize $P$ with a static scale: $\delta_{P}=\frac{1}{448}$ since the original $P$ elements is already in $[0,1]$. We quantize $V$ per-channel to address the channel-wise outliers shown in Fig. 3. Empirical results in Table 7 and Table 16 show the average and the worst accuracy of different data types used for $\widetilde{P},V$ across all layers of CogvideoX. The accumulator is always 32-bit. We can see that the accuracy of E4M3 is very close to that of FP16 and superior to E5M2 and INT8. Most modern GPUs have tensor cores that support FP8 Matmul operations, which are twice as fast as those using FP16.
3.4 FP32 MMA Buffer for FP22 Accumulator
While FP8 quantization for $\tilde{P}V$ above is theoretically accurate in simulation, we observe that the actual CUDA implementation suffers a consistent accuracy degradation. After narrowing down the problem, we find that the accumulator for the mma(f32f8f8f32) instruction on the Ada and Hopper architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are FP8 matrices and $C,D$ are FP32 matrices, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator. When $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the high-precision mma(f16f16f32) instruction. However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32). Consequently, the matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
To mitigate this accuracy loss, we adopt a two-level accumulation strategy, which uses an FP32 buffer to accumulate the values of $\widetilde{P}_{ij}V_{j}$ in FP22. Specifically, we rewrite Eq. (1) as $R_{ij}=\widetilde{P}_{ij}V_{j},O_{ij}=\mathrm{diag}\left(\exp(m_{i,j-1}-m_{ij})\right)O_{i,j-1}+R_{ij}$. Here, two sets of accumulators $R_{ij}$ and $O_{ij}$ are maintained in the register. $R_{ij}$ is computed with the mma(f32f8f8f32) instruction, providing 22 effective bits, which is sufficient since we only accumulate over a small number of $b_{k}$ tokens (e.g., $b_{k}=64$). Then, $R_{ij}$ is accumulated to $O_{ij}$ in the high FP32 precision.
Remark. The two-level accumulation strategy is also implemented in Cutlass (NVIDIA, 2023) and DeepSeekv3 (DeepSeek-AI et al., 2024) for computing activation-weight products in linear layers. To the best of our knowledge, we are the first to discover and investigate the effect of the FP22 accumulator and implement the two-level accumulation for attention.
Optional smooth V technique. We also figure out another way to mitigate the accuracy loss due to the FP22 accumulator when $V$ possesses channel-wise biases: $\overrightarrow{V_{m}}=\mathrm{mean}(V,\text{axis}=0),V=V-\overrightarrow{V_{m}}$. Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$.
Remark. For details on smoothing V, see Appendix A.3. This technique is optional and not employed in our main experiments, as it provides significant benefits only when $V$ exhibits channel-wise bias, which are absent in some models, such as Llama3.1 (see Fig. 3).
Table 2: End-to-end metrics across text, image, and video generation models. ✗ indicates an inability to generate results for evaluation.
Llama3.1 Full-Precision 6.013 0.815 0.635 49.40
HadmdAttn 7.872 0.762 0.500 44.07
SmoothAttn 7.180 0.783 0.541 44.69
SageAttention 6.017 0.812 0.634 49.55
SageAttn2-4b 6.256 0.798 0.607 48.79
SageAttn2-8b 6.019 0.811 0.634 49.59
GLM4 Full-Precision 7.241 0.432 0.743 49.78
HadmdAttn 7.989 0.435 0.669 45.97
SmoothAttn 8.943 0.449 0.592 42.20
SageAttention 7.243 0.433 0.744 49.79
SageAttn2-4b 7.352 0.433 0.725 49.23
SageAttn2-8b 7.242 0.432 0.745 49.60
HadmdAttn 0.1576 0.9933 8.990 2.299
SmoothAttn 0.1559 0.9950 8.812 2.277
SageAttention
FlashAttn3-fp8 0.1562 0.9918 6.531 2.181
SageAttn2-4b 0.1721 0.9978 57.729 52.989 2.884
SageAttn2-8b 0.1775 0.9980 69.492 74.415 2.487
Hunyuan Video Full-Precision 0.1783 0.9995 82.516 75.934 0.604
HadmdAttn 0.1727 0.9989 7.514 0.762 0.175
SmoothAttn 0.1739 0.9988 6.987 0.609 0.148
SageAttention 0.1786 0.9995 82.496 79.843 0.597
FlashAttn3-fp8 0.1742 0.9941 4.433 1.460
SageAttn2-4b 0.1751 0.9995 81.478 65.371 0.610
SageAttn2-8b 0.1782 0.9996 81.786 75.354 0.586
Mochi Full-Precision 0.1798 0.9986 45.549 65.416 1.266
HadmdAttn 0.1733 0.9980 9.053 25.133 0.704
SmoothAttn 0.1687 0.9978 3.383 3.480 0.241
SageAttention 0.1800 0.9987 48.707 63.763 1.269
FlashAttn3-fp8 0.1762 0.9982 14.964 13.711 0.457
SageAttn2-4b 0.1783 0.9986 35.955 43.735 1.137
SageAttn2-8b 0.1797 0.9986 46.760 64.901 1.255
          Flux           Full-Precision          10.960          16.648          26.180          1.009
          HadmdAttn          11.353          18.495          26.123          0.965
          SmoothAttn          11.149          19.017          26.109          0.959
          SageAttention          10.944          16.641          26.171          1.008
          SageAttn2-4b          10.577          17.497          26.141          0.998
          SageAttn2-8b          10.927          16.723          26.175          1.009
                   Stable-Dif          fusion3.5           Full-Precision          14.105          15.646          25.505          0.902
          HadmdAttn          14.259          15.909          25.513          0.886
          SmoothAttn          14.161          15.649          25.510          0.887
          SageAttention          14.140          15.678          25.503          0.902
          SageAttn2-4b          14.097          15.397          25.487          0.895
          SageAttn2-8b          14.106          15.647          25.499          0.901
4 Experiment
Main result. SageAttention2 is faster than FlashAttention2 and xformers by approximately 3x and 4.5x, respectively. Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs and is much more accurate than FlashAttention3(fp8). Our approach maintains end-to-end metrics across language, image, and video generation models.
4.1 Setup
Models. We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation. Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2025), HunyuanVideo (Kong et al., 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
Datasets and metrics. For Details about the datasets and metrics we used, please refer to Appendix. A.6.
Table 3: Two kernel implementations of SageAttention2.
Kernel $\psi_{Q}(Q)$, $\psi_{K}(K)$ $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
SageAttn2-4b INT4 per-thread FP8 per-block and per-channel
SageAttn2-8b INT8 per-thread FP8 per-block and per-channel
Implemetation. We implement two attention kernels as shown in Table 3 using CUDA.
Baselines. (1) SmoothAttn. Following Qserve (Lin et al., 2024), we apply smooth quant for $Q,K$ with smoothing factor $\alpha=0.5$. (2) HadmdAttn. Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for $Q,K$ before INT4 quantization. (3) SageAttention (Zhang et al., 2025), which uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$. (4) FlashAttn3(fp8), the FP8 version of FlashAttention3, which only runs on Hopper GPUs.
Table 4: Average accuracy across all layers of CogvideoX using different smoothing methods.
Method CosSim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
None 80.04% 0.3906 0.2223
HadmdAttn 79.77% 0.3782 0.2180
SmoothAttn 90.21% 0.3383 0.1952
Smooth K 98.07% 0.1493 0.0743
Smooth Q 98.30% 0.1250 0.0712
Smooth Q+K 99.46% 0.0648 0.0334
Table 5: End-to-end metrics comparison, where $Q,K$ are quantized into INT4, while $\widetilde{P},V$ stay in full precision.
Q, K Smooth (Q+K) Llama3.1 (Lambda) $\uparrow$ Llama3.1 (WikiText) $\downarrow$ CogVideoX (vqa-t) $\uparrow$
Full-Precision - 81.5% 6.013 75.360
INT4 Quantization 72.6% 11.698 24.670
80.8% 6.219 75.147